# Are Dense Labels Always Necessary for 3D Object Detection from Point Cloud?

### v1

@@ Abstract
Abstract
In this supplementary material, we firstly present more analysis of visualization comparison, containing different results on the KITTI dataset and the Waymo dataset for the proposed method as well as other ones. Besides, we provide more analysis of pseudo instances to comprehensively explore the influences. Finally, we demonstrate the algorithm of the learning process from streaming unlabeled scenes to ease the reproduction of the proposed method.
@@ 1 Introduction
Autonomous driving, which aims to enable vehicles to perceive the surrounding environments intelligently and move safely with little or no human effort, has attracted much attention recently [16]. As one of the fundamental problems in autonomous driving, detecting 3D objects accurately from a point cloud is crucial for real-world applications. With the advances of deep learning techniques in computer vision and robotics, a number of approaches [14, 23, 15, 12, 28] based on either voxel-wise or point-wise features have been proposed and achieved state-of-the-art (SOTA) performance on large-scale benchmark datasets [13, 9].
@@
Figure 2: Illustration of 3D moderate AP (Average Precision) vs. annotation cost, tested on KITTI  [13] validation set and the AP is calculated with 11 recall positions at IoU 0.7 for the car to compare with the previous methods. Compared with SOTA fully-supervised detectors, our SS3D++ yields promising results with a far lower annotation demand. Besides, when providing less annotations than the remarkably weakly-supervised WS3D [24], our SS3D++ still shows profitable detection performance.
However, most of the proposed 3D object detectors require large-scale, precisely densely-annotated datasets. Unfortunately, collecting high-quality 3D annotations is often time-consuming and labor-intensive compared with 2D image object annotations, in which annotators have to switch viewpoints or zoom in and out throughout a 3D scene carefully for annotating each 3D object, and take hundreds of hours to annotate just one hour of driving data [24]. This poses a huge challenge faced by 3D object detection systems in real-world settings, which requires massive, exhaustively-labeled 3D densely-supervised training data. Therefore, in order to promote the deployment of current 3D object detection systems, it is necessary and significant to reduce the cumbersome data annotation process, and meanwhile achieve 3D object detectors with on-par fully-supervised detection performance.
@@
To mitigate this challenge, some weakly-supervised and semi-supervised approaches [59, 24, 11, 60] have been proposed, which attempt to learn 3D object detection from relatively easily acquired annotations, e.g, leveraging a few high-quality labeled and much unlabeled data [11], center-click point annotation [24] or 2D image-level bounding box annotation [59], as depicted in Fig. 1. Though these methods have substantially reduced the heavy data annotation burden, there exist several deficiencies for applying them in practice. On the one hand, as shown in Fig. 2, their performances have a large margin deterioration compared with the SOTA results of fully-supervised methods. Such dissatisfied performances would not meet the requirements for autonomous driving and practical applications. On the other hand, these low-cost annotations are hardly directly trained using current SOTA fully-supervised detectors, e.g., WS3D [24] need to additionally elaborately designed detector architecture, 3DIoUMatch [11] needs to use additive IoU-estimation module to predict 3D IoU value. These specific detector designs for various weakly-supervised forms relatively restrict the potential to benefit from up-to-date SOTA fully-supervised 3D detectors. Both of the above two aspects may prevent the applicability of these weakly-supervised approaches in real-world settings.
@@
To further promote the deployment of 3D object detection systems, in this paper, we propose a novel sparsely-supervised method that learns 3D object detectors from sparse annotation data in which we just annotate one 3D object per scene, as illustrated in Fig. 1. Intuitively, such novel weakly-supervised annotation form facilitates learning information of unlabeled objects, since infra-scene information transfer is much easier than cross-scene information transfer. Comparatively, semi-supervised scenario [11] may fall into the suboptimal solution due to limited information transfer from labeled to large discrepancy unlabeled scenes, and existing weakly-supervised scenarios may suffer intractable location issues due to significant disparity between 2D-level [59] or point-level [24] weak annotations and 3D-level annotations. Therefore, the proposed sparse annotation strategy has the potential to achieve better performance.
@@
However, such sparsely annotated training data provides inexact and incomplete supervision, naturally raising several new challenges. (1) The missing-annotated instances and the region near those instances may be incorrectly marked as background, which can easily deteriorate the detector under the guidance of incorrect negative samples; (2) To eliminate the negative effect of inexact supervision from positive samples, we are required to generate reliable pseudo-annotated instances (high precision), and find possibly many newly annotated instances (high recall); (3) The sparsely-annotated object in each scene is randomly chosen, while the hardness of learning different objects has great differences. This possibly hinders the knowledge transfer from labeled objects to unlabeled objects, and thus impairs the performance on unlabeled instances.
@@
To address these challenges, we propose a novel and effective method for sparsely annotated 3D object detection, namely SS3D++, which can be applied to any existing off-the-shelf fully-supervised 3D detectors. The main idea of our SS3D++ is to progressively mine positive instances and backgrounds with high confidence using the proposed reliable background mining module and missing-annotated instance mining module, and further use them to generate confident fully-annotated scenes by the GT (ground-truth) Sampling [10] data augmentation strategy. Such a strategy could guarantee that generated pseudo annotations of training instances with high recall provide sufficient knowledge for training the 3D detector. By leveraging the mutual amelioration between high-quality training scene generation and 3D detector training, compared with the 3D detector trained with the fully annotated dataset, our SS3D++ can achieve comparable or even better performance on the KITTI dataset [13] with 5×5$\times$5 × less annotations, and obtain 90% performance with 15×15$\times$15 × less annotations on the Waymo dataset [9].
@@
Our sparsely-supervised framework has three appealing characteristics. First, we can achieve high performance approaching or even surpassing SOTA fully-supervised methods, as presented in Table I and Table III about experimental results of the KITTI [13] and Waymo datasets [9]. However, we only require 5×5$\times$5 × less annotations than fully-supervised methods on KITTI dataset [13], as shown in Fig. 2. Besides, our annotation cost is similar or even less than existing weakly-supervised approaches [59, 24, 11]. Specifically, under the similar number of 3D annotation objects, our annotations are relatively more easily acquired than existing weakly-supervised methods, since it is often easy to just annotate one 3D object per scene. Comparatively, 3DIoUMatch [11] needs to annotate dense 3D objects traversing the entire scene, and WS3D [24] needs to additionally provide center-click point annotations for each 3D object in the entire training scene data. Despite its simplicity in annotation manner, our SS3D++ method is able to achieve substantially better performance, and tends to bridge the performance gap between weakly-supervised learning and strong fully-supervised learning. The detailed experimental results can refer to Fig. 11 and Table IV.
@@
Second, our method is detector-agnostic, which can easily benefit from advanced fully-supervised 3D detectors. As illustrated in Tables I and III, we directly use several existing off-the-shelf advanced fully-supervised 3D detectors to train with sparsely annotated data, resulting in a severe deterioration of the detection performance. While we instantiate our SS3D++ framework as these off-the-shelf fully-supervised 3D detectors without modifying the architectures of the detectors, which consistently improves the performance approaching or even surpassing the fully-supervised results. This property is promising to apply our framework into the autonomous driving system, since it is simple and convenient to adapt to up-to-date 3D detectors and refresh the system to obtain better performance. However, existing weakly-supervised methods [59, 24, 11] elaborately designing specific architectures of detectors to address such imperfect information, relatively hardly benefit from advanced fully-supervised 3D detectors.
@@
Third, our SS3D++ method is hopeful to learn from additional unlabeled training scenes. On the one hand, if we have access to additional unlabeled training scenes in an offline manner, as presented in Fig. 11, we can further boost the performance of a 3D detector by making use of these unlabeled data without additional annotation costs. This implies that our method is capable of generating confident pseudo annotations for easily obtained unlabeled point cloud scenes. On the other hand, if additional unlabeled training scenes arrive in an online manner, we can also continuously improve the performance of the 3D detector by making use of these streaming unlabeled data, as shown in Fig. 15 (c). This property is potentially useful for autonomous driving system, which is required to perceive the changing environment and incrementally improve itself.
@@
This work builds upon our conference paper [52] and expands it in various aspects. Firstly, we introduce the hardness of learning different objects to eliminate the influence of randomly selecting a sparse-annotated object in a scene in Section 4.2.1. Secondly, we rigorously formulate the missing positive instances mining process as a multi-criteria sample selection process, and use sample selection curriculum setups to help to adaptively determine high-confidence positive instances involved in the training pool in Section 4.2. Thirdly, we improve the previous algorithm by leveraging the mutual benefit between 3D detector training and confident fully-annotated scene generation in Section 4.4. Fourthly, extensive ablation studies (see Sections 5.5 and 5.7) and experiments on large-scale Waymo Open Dataset (see Table III) are conducted to demonstrate the effectiveness and universality of our method. Fifthly, we make an in-depth comparison with other weakly-supervised methods in Section 5.4. Finally, we show that the proposed method has the potential for utilizing additional unlabeled training scenes to ameliorate the model in Section 5.6.
@@ 2 Related Work
[…]
@@
Weakly/Semi-supervised 3D object detection. To reduce annotations of 3D objects, the weakly-supervised learning strategy is adopted in WS3D [24], which is achieved by a two-stage architecture based on the click-annotated scheme. WS3D[24] generated cylindrical object proposals by click-annotated scenes in stage-1 and refined the proposals to get cuboids using slight well-labeled instances in stage-2. However, the supervision information provided by the weakly-supervised point annotation is too weak, so that a certain amount of full annotations have to be provided additionally. Meanwhile, some semi-supervised methods [26, 67, 66, 68] leveraged the mutual teacher-student framework to transfer information from labeled scenes to unlabeled scenes. 3DIoUMatch [11] conducted a pioneering work for outdoor scenes, which estimate 3D IoU as a localization metric and set a self-adjusted threshold to filter pseudo labels. Though these methods have substantially reduce the heavy data annotation burden, they have a large margin performance gap with fully-supervised methods, and relatively hardly benefit from advanced 3D detectors. All of these prevent potential practical applicability of 3D object detection. In this paper, we attempt to explore a new weakly-supervised form, sparse supervision, and facilitate 3D object detection close to practical applications.
@@
Curriculum learning for object detection. Curriculum learning [37, 43, 47] is a training strategy that trains a machine learning model in a meaningful order, demonstrating that learning from easy to hard examples can be beneficial. This strategy has been successfully employed in object detection[38, 39, 48, 49, 51, 55, 54]. [38] proposed a curriculum learning strategy to feed training images into the weakly supervised object localization learning loop in an order from images containing bigger objects down to smaller ones. Based on [50], C-SPCL[49] proposed a novel collaborative self-paced curriculum learning framework for weakly supervised object detection, which combined the instance-level confidence inference, the image-level confidence inference, and the self-paced learning mechanism to increase the learning robustness. Under very few annotated samples, [58] embedded multi-modal learning and curriculum learning to ameliorate the trade-off between precision and recall in training data production. Different from these methods tackling 2D object detection, we explore curriculum learning to 3D object detection.
@@
where $\lambda$ is the age parameter for controlling the learning pace, and $\boldsymbol{v}=\left[v_{1},v_{2},\ldots,v_{N}\right]^{\top}\in[0,1]^{N}$ is the latent weight variable. The above learning objective is generally optimized with ACS (Alternative Convex Search) [50]. Concretely, it alternatively optimize $\phi$ and $\boldsymbol{v}$ while fixing the other. With the fixed $\boldsymbol{v}$, the global optimum $\phi^{*}$ can be learned by the existing off-the-shelf supervised learning. Then, with fixed $\phi$, the close-formed optimal solution for $\boldsymbol{v}^{*}$ can be obtained by solving:
$v_{i}^{*}=\left\{\begin{array}[]{lr}1,&l_{i}<\lambda,\\ 0,&\text{ otherwise. }\end{array}\right.$ (4)
@@
Figure 3: The pipeline of our SS3D++ algorithm. We alternatively improve 3D detector training and confident fully-annotated scene generation in a unified learning scheme. To efficiently train the 3D detector, we construct confident fully-annotated scenes based on the missing-annotated instance mining module, the reliable background mining module , and the GT sampling data augmentation strategy. By leveraging the mutual amelioration between high-quality training scene generation and 3D detector training, the obtained 3D detector becomes more robust and well-performing.

### v2

@@ 1 Introduction
Autonomous driving, which aims to enable vehicles to perceive the surrounding environments intelligently, has attracted much attention recently [27]. As one of the fundamental problems in autonomous driving, detecting 3D objects accurately from a point cloud is crucial. With the development of deep learning techniques, a number of approaches [44, 5, 87, 86] based on either voxel-wise or point-wise features have been proposed and achieved state-of-the-art (SOTA) performance on large-scale datasets [10, 57]. However, most of the proposed methods require precisely densely-annotated datasets. Unfortunately, collecting high-quality 3D annotations is often time-consuming and labor-intensive, requiring hundreds of hours to annotate just one hour of driving data [29]. Therefore, it is necessary and significant to reduce the cumbersome annotation process.
@@
Figure 2: Illustration of 3D moderate AP (Average Precision) vs. annotation cost, tested on KITTI  [10] validation set and the AP is calculated with 40 recall positions at IoU 0.7 for the car to compare with the previous methods. Compared with SOTA semi-supervised detectors, our SS3D++ yields promising results with a far lower annotation demand. Besides, when providing less annotations than the remarkably weakly-supervised methods, our SS3D++ still shows profitable detection performance. The cost calculation is based on analysis in FGR [68] and ViT-WSS3D [80].
@@
To mitigate this challenge, some weakly-supervised and semi-supervised approaches [68, 63, 39] have been proposed, which attempt to learn 3D detector from relatively easily acquired annotations, e.g, leveraging a few labeled and much unlabeled data [63], center-click point [29] or 2D image-level bounding box annotation [68], as depicted in Fig. 1. Though these methods have substantially reduced the heavy annotation burden, there exist several deficiencies for practice. On the one hand, their performances are unsatisfied and would not meet the requirements for practical applications. On the other hand, these low-cost annotations are hardly directly trained using fully-supervised detectors. These specific detector designs for weakly-supervised forms relatively restrict the potential to benefit from up-to-date fully-supervised detectors. Both of the above two aspects may prevent the applicability in real-world settings.
@@
To promote the deployment of 3D detector systems, we propose a novel sparsely-supervised method that learns 3D detectors from sparse annotation in which we just annotate one 3D object per scene, as show in Fig. 1. Intuitively, such novel sparse annotation form facilitates learning information of unlabeled objects, since intra-scene information transfer is much easier than inter-scene information transfer. Comparatively, semi-supervised scenario [63] may fall into the suboptimal solution due to limited information transfer from labeled to large discrepancy unlabeled scenes, and existing weakly-supervised methods may suffer intractable location issues due to significant disparity between 2D-level [68] or point-level [29] weak annotations and 3D-level annotations. Therefore, the proposed sparse annotation has the potential to achieve better performance.
@@
However, such sparse annotation data provides inexact and incomplete supervision, naturally raising several new challenges. (1) The missing-annotated instances and the region near those instances may be incorrectly marked as background, which can easily deteriorate the detector under the guidance of incorrect negative samples; (2) To eliminate the negative effect of inexact supervision from positive samples, we are required to generate reliable pseudo-annotated instances (high precision), and find possibly many newly annotated instances (high recall); (3) The sparsely-annotated object in each scene is randomly chosen, while the hardness of learning different objects has great differences. This possibly hinders the knowledge transfer from labeled to unlabeled objects, and impairs the performance.
@@
To address these challenges, we propose an effective method for sparsely annotated 3D object detection, namely SS3D++. The main idea of SS3D++ is to progressively mine positive instances and backgrounds with high confidence using our reliable background mining and missing-annotated instance mining modules, and further use them to generate confident fully-annotated scenes by the GT sampling [74] data augmentation strategy. Such a strategy guarantees that generated pseudo annotations of training scenes with high recall, providing sufficient knowledge.
@@
Our sparsely-supervised framework has three appealing characteristics. First, we can achieve high performance or even surpassing SOTA fully-supervised methods, as presented in Tab. I and Tab. III about experimental results of the KITTI [10] and Waymo datasets [57]. However, we only require $5\times$ less annotations than fully-supervised methods on KITTI dataset [10], as shown in Fig. 2. Besides, when our annotation cost is similar or even less than existing weakly-supervised approaches [68, 29, 63], we can achieve higher results. Specifically, under the similar number of 3D annotation objects, our annotations are relatively more easily acquired than existing weakly-supervised methods, since it is often easy to just annotate one 3D object per scene.
@@
Second, our method is detector-agnostic, which can easily benefit from advanced fully-supervised 3D detectors. As illustrated in Tables I and III, we instantiate our SS3D++ framework as these fully-supervised detectors without modifying the architectures, which consistently improves the performance approaching or even surpassing the fully-supervised results. This property is promising to apply our framework into the autonomous driving system, since it is simple and convenient to adapt to up-to-date 3D detectors and refresh the system to obtain better performance. However, existing weakly-supervised methods [68, 29] elaborately designing specific architectures of detectors to address such imperfect information, relatively hardly benefit from advanced fully-supervised 3D detectors.
@@
Third, our SS3D++ method is hopeful to learn from additional unlabeled training scenes. On the one hand, if we have access to additional unlabeled training scenes in an offline manner, as presented in Fig. 11, we can further boost the performance of a 3D detector by making use of these unlabeled data. This implies that our method is capable of generating confident pseudo annotations for easily obtained unlabeled scenes. On the other hand, if additional unlabeled training scenes arrive in an online manner, we can also continuously improve the performance of the 3D detector by making use of these streaming unlabeled data, as shown in Fig. 14 (c). This property is potentially useful for autonomous driving system, which is required to perceive the changing environment and incrementally improve itself.
@@
This work builds upon our conference paper [23] which has attracted promising attention in the research community. Subsequent studies have further advanced sparse supervision through various approaches, including prototype-based object mining [88], cross-modal prompts [85], hard instance enhancement [73], and contrastive instance feature mining [72]. In contrast, the work in this paper significantly extends the original conference paper in various aspects. Firstly, we introduce the hardness of learning different objects to eliminate the influence of randomly selecting a sparse-annotated object in a scene in Sec. 4.2.1. Secondly, we rigorously formulate the missing positive instances mining as a multi-criteria sample selection process, and use sample selection curriculum setups to adaptively determine high-confidence positive instances involved in the training pool in Sec. 4.2. Thirdly, we improve the algorithm by leveraging the mutual benefit between 3D detector training and confident fully-annotated scene generation in Sec. 4.4. Fourthly, extensive ablation studies (see Sec. 5.5) and experiments on large-scale Waymo dataset (see Tab. III and Tab. VI) are conducted to validate the effectiveness and universality. Fifthly, we make an in-depth comparison with other weakly-supervised methods in Sec. 5.4. Finally, we show that the proposed method has the potential for utilizing additional unlabeled scenes to ameliorate the model in Sec. 5.6.
@@ 2 Related Work
Fully-supervised 3D object detection. The existing 3D detection methods can be broadly categorized into two types: voxel-based methods [74, 87, 86, 16, 11, 61, 83, 15] and point-based methods [36, 76, 75, 33, 46, 82]. Thanks to the seminal PointNet series [34, 37], point-based methods directly take the raw irregular points as input to extract local and global features. PointRCNN [44] directly generated 3D proposals from raw points in a bottom-up manner. Part-A2 [45] further improved the PointRCNN by exploring the rich information in intra-object part locations. For voxel-based methods, voxelization is a common measure for irregular point clouds to apply traditional 2D or 3D convolution. PV-RCNN [42] extended the SECOND [74] by leveraging a keypoints branch to learn discriminative features. PV-RCNN++ [43] proposed the VectorPool aggregation and sectorized keypoints sampling strategy to further enhance the PV-RCNN. Voxel-RCNN [5] achieved a careful balance between performance and efficiency based on the voxel representations. CenterPoint [78] firstly used a keypoint detector to detect centers of objects and then regressed to other attributes. MPPNet [4] leveraged the online tracker to link 3D objects in the temporal dimension. Meanwhile, some methods [35, 26] explored the rich temporal information among frames to facilitate the 3D object detection. SST [8] and DSVT [60] utilized transformers as replacements for sparse CNNs.
@@
Weakly/Semi-supervised 3D object detection. To reduce annotations of 3D objects, the weakly-supervised learning strategy is adopted in WS3D [29], which is achieved by a two-stage architecture based on the click-annotated scheme. WS3D [29] generated cylindrical object proposals by click-annotated scenes in stage-1 and refined the proposals to get cuboids using slight well-labeled instances in stage-2. ViT-WSS3D [80] studied weakly semi-supervised 3D detection with massive click-annotated data and a small number of fully labeled 3D objects. In addition to the weakly-supervised methods based on the point-level annotations, there are also methods [68, 21, 38] that utilize 2D bounding boxes for alleviating the cumbersome 3D annotation process. MTrans [20] proposed a cross-modal Transformer to complement point cloud with dense visual images. Following MTrans, CAT [38] introduced a global encoder to interpret contextual information, thereby removing the visual content. Further, MEDL-U [31] introduced an evidential deep learning based uncertainty estimation framework to address the ambiguities in pseudo labels. Though point-level and 2D-level weakly supervised methods have subtantially reduced the heavy data annotation, their dissatisfied performances and elaborately designed structure may prevent the applicability. Besides, the supervision provided by the weakly-supervised point-level or 2D-level annotation is too weak, so that a certain amount of full annotations have to be provided additionally.
Meanwhile, some semi-supervised methods [84, 32, 77, 17] leveraged the mutual teacher-student framework to transfer information from labeled scenes to unlabeled scenes. 3DIoUMatch [63] conducted a pioneering work for outdoor scenes, which estimate 3D IoU as a localization metric and set a self-adjusted threshold to filter pseudo labels. HSSDA [22] leveraged the hierarchical supervision and shuffle data augmentation to facilitate the semi-supervised learning. A-teacher [62] designed an online asymmetric framework with attention-based refinement to improve the quality of pseudo labels. PPTM [70] generated strong pseudo labels with cross-mixes pillars data augmentation. Though these methods have substantially reduced the heavy data annotation burden, they are relatively hard to benefit from advanced 3D detectors. Besides, the semi-supervised methods need to annotated all the 3D objects within a scene, in which annotators have to switch viewpoints or zoom in and out throughout a 3D scene carefully. All of these prevent potential practical applicability of 3D object detection. In this paper, we attempt to explore a new weakly-supervised form, sparse supervision, and facilitate 3D object detection close to practical applications. The proposed sparse annotation, which requires labeling only a subset of objects, significantly reduces the number of point clouds that need attention. This reduction not only lessens the burden on annotators but also minimizes the potential for errors during the annotation process, thereby greatly lowering the costs associated with manual work.
@@
Curriculum learning for object detection. Due to that a part of instances are missing-annotated, weight updated of the network may be misguided significantly when gradients back-propagated. To address this issue, existing 2D detection methods employed re-weight or re-calibrates strategy on the loss of RoIs (regions of interest) to eliminate the effect of unlabeled instances. Soft sampling [71] utilized overlaps between RoIs and annotated instances to re-weight the loss. Background recalibration loss[81] based on focal loss[19] regarded the unlabeled instances as hard-negative samples and re-calibrates their losses, which is only applicable to single-stage detectors. Especially, part-aware sampling[30] ignored the classification loss for part categories by using human intuition for the hierarchical relation between labeled and unlabeled instances. Co-mining [65] proposed a co-generation module to convert the unlabeled instances as positive supervisions. However, these sparsely annotated object detection methods are all for 2D image objects. Due to the modal difference between 2D images and 3D point cloud, these methods can not be directly applied to our 3D object detection task. For example, in KITTI [10], 3D objects are naturally separated, which means the overlaps among objects are zero and the hierarchical relation between objects does not exist. Compare with the re-weight and re-calibrates methods, in this paper, we propose a novel method for sparsely annotated 3D object detection which leverages a simple but effective background mining module and a missing-annotated instance mining module to mine confident positive instances and backgrounds, which is key for training detectors with high performance.
Curriculum learning for object detection. Curriculum learning [1, 66, 55] is a training strategy that trains a machine learning model in a meaningful order, demonstrating that learning from easy to hard examples can be beneficial. This strategy has been successfully employed in object detection[41, 64, 47, 79, 18, 56, 40]. [41] proposed a curriculum learning strategy to feed training images into the weakly supervised object localization learning loop in an order from images containing bigger objects down to smaller ones. Based on [12], C-SPCL[79] proposed a novel collaborative self-paced curriculum learning framework for weakly supervised object detection, which combined the instance-level confidence inference, the image-level confidence inference, and the self-paced learning mechanism to increase the learning robustness. Under very few annotated samples, [6] embedded multi-modal learning and curriculum learning to ameliorate the trade-off between precision and recall in training data production. Different from these methods tackling 2D detection, we explore curriculum learning to 3D detection.
@@
where $\lambda$ is the age parameter for controlling the learning pace, and $\boldsymbol{v}=\left[v_{1},v_{2},\ldots,v_{N}\right]^{\top}\in[0,1]^{N}$ is the latent weight variable. The above learning objective is generally optimized with ACS (Alternative Convex Search) [12]. Concretely, it alternatively optimizes $\phi$ and $\boldsymbol{v}$ while fixing the other. With the fixed $\boldsymbol{v}$, the global optimum $\phi^{*}$ can be learned by the existing off-the-shelf supervised learning. Then, with fixed $\phi$, the close-formed solution for $\boldsymbol{v}^{*}$ can be obtained by solving:
[…]
@@
Figure 3: The pipeline of our SS3D++ algorithm. We alternatively improve 3D detector training and confident fully-annotated scene generation in a unified learning scheme. To efficiently train the 3D detector, we construct confident fully-annotated scenes based on the missing-annotated instance mining module, the reliable background mining module , and the GT sampling data augmentation strategy. By leveraging the mutual amelioration between high-quality training scene generation and 3D detector training, the obtained 3D detector becomes more robust.